# Data-Juicer 2.0: Cloud-Scale Adaptive Data Processing for and with Foundation Models

### v2

@@ 4.3 Processing Job Control
End-to-End Workflow Orchestration. Our Executor module integrates configurable data pipelines with monitoring/checkpointing capabilities, codified into reusable data recipes (the red box in Fig. 1). We also provide a sandbox suite that enables data-model co-development through template workflows connecting to model training/evaluation infrastructures chen2024djsandbox , allowing cost-effective exploration of data-compute effects before full deployment. More implementation details are in Appendix E.4.
@@ 5.4 The Case of Large Scales
Figure 4: Processing time comparison across computing engines, dataset sizes, and workloads.
@@
Additionally, the adaptive subset splitting capability for Ray mode, introduced in Sec. 4.1, offers 2$\sim$3x acceleration in our experiments. For example, with 100 Ray-DLC nodes and 12,800 cores, our subset splitting method reduces the processing time of the 12,500x dataset from over 5,000s to about 2,000s. Analysis of Ray’s resource utilization shows that its automatic block-splitting strategy incurs high network communication costs and low CPU utilization when many nodes handle a few dataset files. Our strategy pre-splits datasets according to size and node count, optimizing alignment with Ray’s features and enhancing efficiency. As seen in Fig. 4f, this strategy reduces peak network I/O from about 160MB/s to 60MB/s and achieves more consistent CPU utilization across all nodes.
@@
{tblr}
width = colspec = Q[140]Q[270]Q[240]Q[240], hline1,4 = -0.08em, hline2 = -, # CPU & 200GB Time 1TB Time 5TB Time
@@
Large-scale Deduplication. We tested the MinHash-based RayDeduplicator with datasets from Redpajama together2023redpajama and Common Crawl chunks commoncrawl , sized at 200GB, 1TB, and 5TB, using CPU counts ranging from 640 to 1,280 cores. As shown in Table 1, RayDeduplicator efficiently scales with increased data size and computing resources, indicating its capability to handle large-scale deduplication tasks effectively. When the data size increases by 5x, the processing time increases by 4.02$\sim$5.62x. When the number of CPU cores doubles, the processing time decreases to 58.9$\sim$67.1of the original time. Notably, RayDeduplicator can process 5TB of data in 2.8 hours using 8 Ray nodes (8*160 CPU cores). As a comparison, NVIDIA’s NeMo Curator, which leverages cuDF cuDF and Dask dask , takes 1.8 hours to process 1.1TB of Redpajama data using 64 A100 GPUs (64*6,192 CUDA cores), as shown in nemo-dedup-results .
@@
5.5 More Ablation Studies on Runtime Adaptations
We conducted several experiments to evaluate the efficiency of enhancements from Data-Juicer core runtime. Overall, OP reordering along with OP fusion can save at most 70.22% processing time for some complex data recipes; Automatic GPU resource allocation helps to prevent OOM errors and saves at most 99% time when processing multimodal datasets with foundation models; Data processing with a batch size of 1000 reduces processing time by at most 84% in most scenarios; And OP-wise hierarchical parallelism speeds up the data processing consistently across all parallel levels (multiprocessing, batched processing, multithreading). One take-away message is that adaptive batching and resource allocation (e.g, via our hierarchical parallelism) can effectively mitigate hardware underutilization caused by irregular computation graphs of data pipelines for foundation models. More details about these ablation studies are presented in Appendix H.2.
6 Related Work
Data Infrastructure for Foundation Models. Modern data processing systems for foundation models need to reconcile conflicting demands between multimodal semantics and computational efficiency wang2023data ; bai2024survey . While existing frameworks excel at structured data analysis apache-spark or text-centric processing DolmaToolkit ; penedo2024datatrove , they struggle with three key gaps. First, current systems lack native support for cross-modal alignment - a critical requirement for quality-sensitive tasks like visual-language pretraining liu2023llava and multimodal synthesis podell2023sdxl . Second, traditional SQL-based optimizations become inadequate for pipelines dominated by neural network inference and fuzzy deduplication schuhmann2021laion , particularly when processing large unstructured records ikegwu2022big . Third, Python-centric workflows of foundation model practitioners Paszke-NeurIPS-2019-Pytorch conflict with JVM-based architectures of legacy big data systems.
Our Position. Next-generation data processing systems for foundation models require three key capabilities to bridge existing gaps: (1) Standard multimodal OPs that handle semantic relationships across vision, language, and audio modalities; (2) Hybrid execution strategies balancing local GPU acceleration with distributed processing for neural data transformations; (3) Python-native interfaces that abstract engine-specific complexities while maintaining compatibility with major computational backends (Hugging Face, Ray, Spark). As detailed in Appendix A, the proposed architecture addresses the tension between semantic-aware processing and scalable execution - a fundamental challenge when handling terabyte-scale multimodal datasets with heterogeneous compute requirements. Our implementation demonstrates how careful system design can bridge the gap between academic prototyping needs and industrial-scale data pipelines.
Difference to Prior Works. As user demands and applications evolved rapidly, the complexity of Data-Juicer 1.0 increased correspondingly. The need for heterogeneous data processing for more modalities with enhanced usability calls for an updated version. Besides, it showed shortcomings in programmability, such as fragmented development workflows across diverse underlying data frame frameworks. Unlike Data-Juicer 1.0 and other prior modality-specific and task-specific toolkits for or with large models Jennings_NeMo-Curator_a_toolkit ; lhoest-etal-2021-datasets ; curator ; penedo2024datatrove , we introduce a new composable OP system that generalizes across data types and training tasks while maintaining scalability optimizations. This contrasts with conventional big data systems ikegwu2022big that prioritize relational algebra over neural network integration, by emphasizing computational efficiency alongside semantic preservation.
@@ 7 Conclusions, Limitations & Future Works
There are several limitations in Data-Juicer 2.0. From the perspective of the computing system, future work includes further enhancing processing adaptability through model-driven agents gao2024agentscope , scalability by optimizing the single-node transmission bottleneck in Ray’s head node ray-head-worker , and efficiency by supporting GPU backend engines like NeMo Curator Jennings_NeMo-Curator_a_toolkit . From the perspective of societal impacts, we now introduce several operators for removing NSFW content, and most operators can only process data in English/Chinese. Expanding processing capabilities to cover multilingualism, data privacy, data safety, and so on, is one of the most important enhancement directions.
@@
[…]
1Table of Contents
Appendix A Discussions on System Challenges and Design Goals
A.1 Data Processing for Foundation Models
Foundation models, especially those beyond text, demands innovative data processing paradigms to address the complexity and diversity of large-scale training datasets [54, 79, 74, 5]. Despite existing systems for big data or text-only models [78, 21, 58, 32], there is a notable lack of open-source frameworks customized for foundation models due to unique design and implementation challenges:
Functionality Challenges from Data Types and Semantics. Data processing frameworks for foundation models primarily target semi-structured textual and multimodal datasets (e.g., JSONL, Parquet, MP4), often disregarding the strongly structured data types prominent in previous big data systems (e.g., database tables) [36, 32]. Significantly, the performance of foundation models depends on the quality, diversity, utility, and safety of data, necessitating a profound semantic understanding and cross-modal alignment to transform raw data into high-quality inputs, thereby enhancing model efficacy [40, 38, 70].
Efficiency Challenges from Data Operations and Pipelines. Conventional engines support a broad range of SQL queries for complex data operations [78, 60]. However, in the context of foundation models, the requirement for traditional grouping and aggregation operations is relatively minimal [5]. Instead, emphasis is placed on simpler operations such as Mapper and Filter, which rely heavily on deep learning models via local inference [57, 11, 73] or API calls [40, 43, 41]. Furthermore, individual records in these datasets can be exceedingly large, surpassing the capabilities of traditional big data engines [47, 31]. When combining multiple data operations into pipelines, it is crucial to maintain efficient data and transformation flows throughout the data lifecycle to prevent bottlenecks stemming from computational overhead, latency, and data integrity.
Usage Challenges from User Interfaces. Traditional big data systems typically use programming interfaces based on Java and Scala, supplemented by SQL. However, practitioners focusing on foundation models predominantly operate within the Python ecosystem, supporting both upstream and downstream repositories and workflows [49, 1, 55]. This shift necessitates designing user interfaces that specifically cater to the new programming needs of foundation model practitioners.
A.2 System Design Goals in Data-Juicer 2.0
Built upon Data-Juicer 1.0, we aim to address challenges identified in Appendix A.1 with the focus on the following new objectives:
Comprehensive Multimodal Processing Abilities. The first basic goal is to support extensive multimodal processing capabilities across diverse domains, such as AI-generated content, spatial intelligence, education, and entertainment. This involves meticulous processes such as collection, cleansing, annotation, and synthesis, which integrate perceptual and cognitive information across modalities such as video, image, text, and audio [10, 52, 2, 22].
Efficient Data Processing Operators and Data Flow. Comprehensive semantic operations are expected through advanced procedures that may be high-cost, such as deep learning model invocations, fuzzy deduplication, and clustering. For basic, high-frequency operations, we need to employ dedicated code optimization and speedup techniques like GPU computation or Cython [6]. Additionally, optimizing the loading, parsing, and storage of multimodal data, while minimizing overhead associated with linking, read/write operations, and network transmissions, is crucial for streamlining processing and enhancing efficiency.
Easy-to-Use Interface. To tackle usage challenges, Data-Juicer 2.0 pursues an intuitive, engine-agnostic abstraction of computational primitives, supporting Python, layered APIs, graphical user interfaces, and even natural language interactions. This design shields users from the inherent complexities of various computational engine implementations and manages the cognitive and usage costs to remain consistent, even as functionalities expand.
Adaptive and Scalable Processing. Last but not least, Data-Juicer 2.0 is designed to dynamically adapt to diverse computational environments and workloads, optimizing performance across frameworks including Hugging Face Dataset, Ray Dataset, and Alibaba Cloud MaxCompute, while considering their respective use cases. For example, OPs based on Hugging Face Dataset align well with large model community compatibility but require enhancements for scalable and distributed processing. Ray-based OPs offer excellent Python compatibility with low interaction overhead, yet they may face single-node bottlenecks for large-scale clusters involving data shuffles, such as those required by join operation. OPs based on Spark and MaxCompute necessitate careful resource management and schema considerations due to serialization between Python and the backend engines (e.g., JVM), rendering them less suitable for high-frequency computations.
In the subsequent sections, we will explore the design specifics and implementation of Data-Juicer 2.0 to illustrate how these objectives are achieved.
These design goals guide the development of robust operator functionalities and system implementation, tailored to meet the specific requirements of foundation model data processing. In the subsequent sections, we will explore the design specifics and implementation of Data-Juicer 2.0 to illustrate how these objectives are achieved.
Appendix B List of New Operators in Data-Juicer 2.0
@@ A.1 Data Processing for Foundation Models
Appendix C Showcase of Typical Multimodal Operators
@@
Appendix D User Interfaces
D.1 Low-level APIs
The framework’s core capabilities are exposed through Python-based programmatic interfaces, providing object-oriented logical encapsulation for both the fundamental Data-Juicer-Dataset, Data-Juicer-Operators and other runtime modules. This design delivers developers maximum flexibility and customizability. Processing workflows can be automatically chained by passing a series of OP instances to a loaded Data-Juicer-Dataset object (e.g., data.process([op1, op2])), enabling various operations to be performed on the dataset in a single pass. Additionally, the framework supports applying an instantiated Data-Juicer-Operators to a target dataset (e.g., op.run(data)), enhancing the reusability of OP instances. More details about the Data-Juicer-Dataset and Data-Juicer-Operators are introduced in Sec. 4 and Appendix E. This dual approach—chained processing and individual OP application—optimizes both efficiency and modularity in data processing tasks, while leveraging the inherent advantages of Python’s ecosystem.
D.2 RESTful APIs
@@ D.3 Web & CMD tools
Thanks to the low-level and RESTful APIs, we further establish service-level capabilities across various scenarios in Data-Juicer 2.0, encapsulating scenario-specific processing solutions into built-in data recipes, described as YAML configuration files for end-to-end workflows (an example for video-data-synthesis is shown in Fig. 6). Centered around these data recipes, we provide user-friendly, highly encapsulated command-line interfaces as illustrated in Listing D.3, supporting automated lightweight installation and recipe-level invocation. These tools automatically gather and install the requisite of OPs within the recipe, perform automated OP fusion, and maximize adaptive utilization of computing resources while minimizing user cognitive load and operational cost.
{listing}
[!ht] {minted}[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, linenos, breaklines=true, bgcolor=lightgray!10, rulecolor=black, numbersep=5pt, ]bash dj-install –config recipe.yaml # install dependencies only for this recipe dj-analyze –auto –dataset_path xx.jsonl # default OPs, or via specific recipe dj-process –config recipe.yaml # for both standalone or distributed Representative tools available in Data-Juicer 2.0.
@@
D.4 Natural Language Interaction
@@
[…]
@@
Appendix E Core Runtime Implementation Details
E.1 Execution Mode Configuration
@@
{listing}
[!h] {minted} [ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, linenos, breaklines=true, bgcolor=lightgray!10, rulecolor=black, numbersep=5pt, ]python # Differernt input sources, such as from local files, dataset hubs from Hugging Face and Modelscope, and special types like arXiv, wiki, CommonCrawl, etc. data = DJDataset.load(src_file)
data = data.process(op1) # run a single operator data = data.process(op1).process(op2) # run multiple operators continuously data = data.process([op1, op2]) # run a list of operators
data.export(tgt_file) # tgt_file can be either local path or remote path Some basic interfaces of Data-Juicer-Dataset.
Moreover, following the principle of Template Method Pattern, the Data-Juicer-Dataset class establishes templated workflows for OP executing on datasets and provides several basic and unified interfaces, catering to users with diverse programming preferences, as demonstrated in Listing E.1. The complexities of underlying computational engines and cumbersome runtime issues related to instantiated dataset objects are abstracted away from these interfaces, ensuring transparency and ease of use for end-users and developers alike.
New functionalities can be added by merely implementing new OP classes without altering the internal logic of Data-Juicer-Dataset. The standalone and distributed execution modes automatically recognize and switch configurations based on runtime settings. Coupled with unified function signatures, the data processing can be consolidated into a batched processing mode, not only standardizing the data interface for OP processing but also facilitating the implementation of a robust sample-level fault tolerance mechanism. The internal adaptation features are elaborated in Appendix F.2.
E.2 Data Schema Implementation
{listing}
[!h] {minted} [ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, linenos, breaklines=true, bgcolor=lightgray!10, rulecolor=black, numbersep=5pt, ] json // >>> core contents: texts, dialogs, … // for general pretraining "text": "<__dj__image> desc of image 1 <|__dj__eoc|> desc of image 2 <__dj__image> <|__dj__eoc|>", // for post-tuning "query": "<query2>", "response": "<response2>", "history": [["<query1>", "<response1>"]], // <<< core contents
// $>$ extra data contents: multimodal data paths, … "images": [ "path/to/the/image1", "path/to/the/image2" ], // $<$ extra data contents
// $>$ meta infos and stats, which could be primitive or produced by Data-Juicer "meta": "src": "customized", "version": "0.1" , "stats": "lang": "en", "image_widths": [224, 336] , // $<$ meta infos and stats Illustration of data schema of Data-Juicer 2.0.
@@ E.2 Data Schema Implementation
An example of a data sample is illustrated in Listing E.2. In the data schema, each dataset encapsulates column-stored samples represented by a non-recursive dictionary whose level-1 fields include three categories: (1) the core contents within customizable fields, including “text” that is usually used in the pretraining task, and “query”, “response”, “history” fields represent dialogs in post-tuning tasks, which are directly related to the pretraining or post-tuning procedures in the downstream usage of the dataset; (2) the extra data contents contain the path lists of multimodal data for multimodal datasets; (3) the “meta” information concerning this sample that stems from either its raw data or Data-Juicer’s Mappers for tagging, and the “stats” information calculated by Data-Juicer’s Filters (all Filter OPs consist of a compute_stats() method, utilized for subsequent analysis and conditional filtering).
For multimodal datasets, the content is primarily centered on the textual modality within the “text” field, capable of optionally including several text chunks split by customizable end-of-chunk tokens, which are <|__dj__eoc|> in default. Each chunk serves as a semantic unit, where the associated multimodal data within the same chunk relates to the same topic, thereby aligning with one another. Multimodal data (excluding text) are denoted by ordered customized special tokens in the text (e.g., <__dj__image>), storing entities at designated files, accessible through indexed file paths within the modality fields of samples to facilitate object sharing and elimination of redundant computing. For instance, in the example shown in Listing E.2, the first chunk in the “text” field split by the <__dj__eoc> token contains the first <__dj__image> token and a description to this image, which corresponds to the first image stored in the “images” field.
@@
E.3 Operator Factory & Taxonomy
E.3.1 OP Design Principles

### v3

@@ 5.3 Processing Job Control
End-to-End Workflow Orchestration. Our Executor module integrates configurable data pipelines with monitoring/checkpointing capabilities, codified into reusable data recipes (the red box in Fig. 1). We also provide a sandbox suite that enables data-model co-development through template workflows connecting to model training/evaluation infrastructures chen2024djsandbox , allowing cost-effective exploration of data-compute effects before full deployment. More implementation details are in Appendix D.4. processing solution generation by foundation-model-based agents.
@@ 6.1 Experiment Setup
6.2 Overview of System Performance Gains
Macro-level Scalability We systematically evaluated end-to-end performance by scaling datasets from 1x to 12,500x. The results, presented later from §6.3 to §6.5, confirm the robust scalability of Data-Juicer and provide actionable insights for choosing the optimal compute engine across different operational scales.
Micro-level Optimizations The strong end-to-end performance is underpinned by a suite of targeted optimizations. Below, we highlight the most impactful ones and point to where their effectiveness is validated (more details are in presented in Appendix H.2):
  • Resource Utilization & Adaptive Splitting: Our adaptive data splitting for Ray offers a 2x-3x acceleration on large datasets. The mechanism’s impact on reducing network I/O and improving CPU consistency is visually analyzed in Fig. 4f and discussed in §6.5.
  • Workload-aware OP Reordering: For complex recipes, this strategy, along with OP fusion, can cut processing time by up to 70.22%. The quantitative benefits are detailed in the ablation studies in Fig. 13.
  • Automatic GPU Resource Allocation: Critical for multimodal workloads, this prevents OOM errors and can save up to 99% of processing time. Its performance across various VRAMs is quantified in Table 3.
  • Batched Data Processing: Optimizing batch size and concurrency is key. As shown in Fig. 14a, this can reduce processing time by up to 84%.
A key takeaway is that adaptive mechanisms for batching, resource allocation, and execution planning are crucial for mitigating hardware underutilization in modern data-centric AI pipelines.
@@ 6.5 The Case of Large Scales
Figure 4: Processing time comparison across computing engines, dataset sizes, and workloads. A high-resolution version of subfigure (f) can be found in Appendix, Figure. 20
@@
[…]
@@
# CPU 200GB Time 1TB Time $5TBTime$
@@
[…]
@@ 7 Conclusions, Limitations & Future Works
There are several limitations in Data-Juicer 2.0. From the perspective of the computing system, future work includes further enhancing processing adaptability through model-driven agents gao2024agentscope , scalability by optimizing the single-node transmission bottleneck in Ray’s head node ray-head-worker , and efficiency by supporting GPU backend engines like NeMo Curator Jennings_NeMo-Curator_a_toolkit . Looking ahead, one key direction for future development is enhancing the framework to cover multilingualism, more scenarios (e.g., AI4science, self-driving, embodied intelligence), and a wider range of data governance and safety considerations, such as ensuring data privacy across enterprise-level security protocols. Besides, scaling to the next order of magnitude of processed data is important and requires forward-thinking architectural design like more advanced pipeline optimization, as we discussed in Appendix G.
@@
Appendix A List of New Operators in Data-Juicer 2.0
@@
Appendix B Showcase of Typical Multimodal Operators
@@
Appendix C User Interfaces
C.1 Low-level APIs
The framework’s core capabilities are exposed through Python-based programmatic interfaces, providing object-oriented logical encapsulation for both the fundamental Data-Juicer-Dataset, Data-Juicer-Operators and other runtime modules. This design delivers developers maximum flexibility and customizability. Processing workflows can be automatically chained by passing a series of OP instances to a loaded Data-Juicer-Dataset object (e.g., data.process([op1, op2])), enabling various operations to be performed on the dataset in a single pass. Additionally, the framework supports applying an instantiated Data-Juicer-Operators to a target dataset (e.g., op.run(data)), enhancing the reusability of OP instances. More details about the Data-Juicer-Dataset and Data-Juicer-Operators are introduced in Sec. 5 and Appendix D. This dual approach—chained processing and individual OP application—optimizes both efficiency and modularity in data processing tasks, while leveraging the inherent advantages of Python’s ecosystem.
C.2 RESTful APIs
@@ C.3 Web & CMD tools
Thanks to the low-level and RESTful APIs, we further establish service-level capabilities across various scenarios in Data-Juicer 2.0, encapsulating scenario-specific processing solutions into built-in data recipes, described as YAML configuration files for end-to-end workflows (an example for video-data-synthesis is shown in Fig. 6). Centered around these data recipes, we provide user-friendly, highly encapsulated command-line interfaces as illustrated in Listing 7, supporting automated lightweight installation and recipe-level invocation. These tools automatically gather and install the requisite of OPs within the recipe, perform automated OP fusion, and maximize adaptive utilization of computing resources while minimizing user cognitive load and operational cost.
dj-install config recipe.yaml # install dependencies only for this recipe
dj-analyze auto dataset_path $x$.jsonl # default OPs, or via specific recipe
dj-process config recipe.yaml # for both standalone or distributed
Figure 7: Representative tools available in Data-Juicer 2.0.
@@
C.4 Natural Language Interaction
@@
Assistant: The analyzed statistics are displayed in the image below: [Uncaptioned image] Assistant: Based on the analysis, please enter the minimum and maximum threshold values for filtering. User: 80 120 System: Execute function run_text_length_filter
@@
Appendix D Core Runtime Implementation Details
D.1 Execution Mode Configuration
@@
# Differernt input sources, such as from local files, dataset hubs from Hugging Face and Modelscope, and special types like arXiv, wiki, CommonCrawl, etc.
data = DJDataset.load(src_file)
[…]
data = data.process(op1).process(op2) # run multiple operators continuously
data = data.process([op1, op2]) # run a list of operators
[…]
Figure 8: Some basic interfaces of Data-Juicer-Dataset.
Moreover, following the principle of Template Method Pattern, the Data-Juicer-Dataset class establishes templated workflows for OP executing on datasets and provides several basic and unified interfaces, catering to users with diverse programming preferences, as demonstrated in Listing 8. The complexities of underlying computational engines and cumbersome runtime issues related to instantiated dataset objects are abstracted away from these interfaces, ensuring transparency and ease of use for end-users and developers alike.
New functionalities can be added by merely implementing new OP classes without altering the internal logic of Data-Juicer-Dataset. The standalone and distributed execution modes automatically recognize and switch configurations based on runtime settings. Coupled with unified function signatures, the data processing can be consolidated into a batched processing mode, not only standardizing the data interface for OP processing but also facilitating the implementation of a robust sample-level fault tolerance mechanism. The internal adaptation features are elaborated in Appendix E.2.
D.2 Data Schema Implementation
// >>> core contents: texts, dialogs,
// for general pretraining
"text": "<__dj__image> desc of image 1 <|__dj__eoc|> desc of image 2 <__dj__image> <|__dj__eoc|>",
// for post-tuning
"query": "<query2>",
"response": "<response2>",
"history": [["<query1>", "<response1>"]],
// <<< core contents
[…]
"images": [
"path/to/the/image1",
"path/to/the/image2"
// <<< extra data contents
[…]
"meta": {
"src": "customized",
"version": "0.1"
"stats": {
"lang": "en",
"image_widths": [224, 336]
// <<< meta infos and stats
Figure 9: Illustration of data schema of Data-Juicer 2.0.
@@ D.2 Data Schema Implementation
An example of a data sample is illustrated in Listing 9. In the data schema, each dataset encapsulates column-stored samples represented by a non-recursive dictionary whose level-1 fields include three categories: (1) the core contents within customizable fields, including “text” that is usually used in the pretraining task, and “query”, “response”, “history” fields represent dialogs in post-tuning tasks, which are directly related to the pretraining or post-tuning procedures in the downstream usage of the dataset; (2) the extra data contents contain the path lists of multimodal data for multimodal datasets; (3) the “meta” information concerning this sample that stems from either its raw data or Data-Juicer’s Mappers for tagging, and the “stats” information calculated by Data-Juicer’s Filters (all Filter OPs consist of a compute_stats() method, utilized for subsequent analysis and conditional filtering).
For multimodal datasets, the content is primarily centered on the textual modality within the “text” field, capable of optionally including several text chunks split by customizable end-of-chunk tokens, which are <|__dj__eoc|> in default. Each chunk serves as a semantic unit, where the associated multimodal data within the same chunk relates to the same topic, thereby aligning with one another. Multimodal data (excluding text) are denoted by ordered customized special tokens in the text (e.g., <__dj__image>), storing entities at designated files, accessible through indexed file paths within the modality fields of samples to facilitate object sharing and elimination of redundant computing. For instance, in the example shown in Listing 9, the first chunk in the “text” field split by the <__dj__eoc> token contains the first <__dj__image> token and a description to this image, which corresponds to the first image stored in the “images” field.
@@
D.3 Operator Factory & Taxonomy
D.3.1 OP Design Principles